OpenStudy (anonymous):

solve and check 3d^2+14d-5=0

OpenStudy (anonymous):

d = -5,1/3

OpenStudy (anonymous):

did you try factoring? if not, use quadratic formula

OpenStudy (anonymous):

do i have to show

OpenStudy (anonymous):

factor 15-1

OpenStudy (anonymous):

hint:now that problemsolver gave you the answer, you know how to factor it

OpenStudy (anonymous):

please show work if you can I want to fully understand

OpenStudy (anonymous):

\[3d ^{2} +14d -5 = (3d -1 )(d+5)\] \[(3d-1) = 0\] \[3d = 1\] \[d = 1/3\] \[d + 5 = 0\] \[d\]

OpenStudy (anonymous):

you want to write it as \[(3d+\text{something})(d+\text{something else})=0\]

OpenStudy (anonymous):

d= -5

OpenStudy (anonymous):

Not sure why it truncated.

OpenStudy (anonymous):

and since you have the answer as \[-5,\frac{1}{3}\] you know it must factor as \[(d+5)(3d-1)=0\]

OpenStudy (anonymous):

thank u....